# SimuHome: A Temporal- and Environment-Aware Benchmark for Smart Home LLM Agents

### v1

@@ 7 Conclusion
We propose SimuHome, a Matter-aligned simulator and benchmark that reproducibly evaluates smart home LLM agents under realistic, dynamically changing conditions. We model 4 environmental variables (i.e., temperature, illuminance, humidity, air quality) and 17 device types with time-based effects and strict reproducibility, enabling near drop-in transfer to real Matter-compliant devices. We provide 600 episodes across 12 query types with feasible and infeasible variants, packaging each episode with an initial state, a verifiable goal, a natural-language query, and required actions for process-aware, objective scoring. We score feasible tasks by final state-to-goal comparison in the simulator and assess infeasible logic checks with LLM judge that shows high agreement with human evaluation. We evaluate 11 agents under the ReAct setup. Current open- and closed-source LLMs handle explicit control and simple retrieval well but often fail at latent-intent inference, live-state verification, and temporal scheduling; GPT-4.1 achieves an overall success rate of 54% across all query types.
Appendix A Appendix
A.1 Infeasible Query Types
QT1 Infeasible.
@@ A.1 Infeasible Query Types
This evaluates the ability to identify requests based on a false premise, such as asking for information about non-existent devices or unsupported attributes. For example, for the request “Can you tell me the vendor ID for the air purifier in the living room?”, the agent must check the list of devices in the living room, confirm the absence of an air purifier, and explain that the request’s premise is invalid.
QT2 Infeasible.
@@
This assesses the ability to identify situations where, even if the user’s intent is correctly inferred, the goal is impossible to achieve due to environmental constraints or device limitations. For example, in response to “The living room feels like a sauna”, the agent must verify that the living room’s cooling system is already operating at maximum capacity and explain, with supporting reasons, why further cooling is not possible.
QT3 Infeasible.
@@
This evaluates the ability to identify and reject a command to control a non-existent device. For example, for the request “Turn on the humidifier in the living room”, the agent must check the device list for the living room and confirm the absence of a humidifier. It should then explain that the request cannot be fulfilled and terminate the task without altering any device’s state.
QT4-1 Infeasible.
@@
This assesses the ability to identify and explain situations where a scheduling request is invalid because the user’s specified relative and absolute times are contradictory, or because the user has a misunderstanding of the current time. For example, if a user asks, “It’s 6 p.m. now, right? Turn on the kitchen light five minutes later at 6:05 p.m.”, but the actual time is not 6 p.m., the agent must check the current time, detect the discrepancy between the relative expression “five minutes later” and the absolute time “6:05 p.m.”, and clearly explain the contradiction.
QT4-2 Infeasible.
@@
This evaluates the ability to identify and explain, with evidence, requests where the user incorrectly assumes a device’s completion time or creates a contradiction by providing both relative and absolute times. For example, suppose a washer is set to finish at 6:30 p.m., but the user requests, “I think the washer finishes at 6 p.m., so start the dehumidifier at 5:50 p.m., which is 10 minutes before it finishes”. The agent must check the washer’s actual estimated completion time. It then needs to point out that the user’s assumption (6 p.m.), the relative expression (“10 minutes before”), and the absolute time (“5:50 p.m.”) are all inconsistent. The agent must not register the schedule until the contradiction is resolved and should ask the user to reconfirm the correct timing.
QT4-3 Infeasible.
@@
This evaluates the ability to identify and explain that a requested deadline is physically impossible to meet, given the current progress of two operating devices. For example, if the user requests, “Guests arrive at 6 p.m., so ensure both the washer and the dishwasher are completed by 5:30 p.m.”, the agent must check the current time and the minimum time required for each device to finish. Based on this, it should explain with clear reasoning why a 5:30 p.m. completion is not feasible and suggest the earliest possible completion time or an alternative sequential plan.
A.2 List of Tools
Table 3: Tool List for Agent
@@ A.2 List of Tools
finish Complete the task and return the final natural-language answer. answer (str, required): Final response text.
execute_command Execute a command on a device (e.g., turn on light, set level, set setpoint). device_id (str, req); endpoint_id (int, req); cluster_id (str, req); command_id (str, req); args (dict, req).
write_attribute Directly set a device attribute value. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req); value (any, req).
get_all_attributes Get all attributes of a device. device_id (str, req).
get_attribute Get a specific attribute of a device. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req).
get_device_structure Get device structure (endpoints, clusters, attributes, and commands). device_id (str, req).
get_rooms Get all rooms in the home along with their display names. (none)
get_room_devices Get all devices in a room. room_id (str, req).
get_room_states Get environmental states of a room (temperature, humidity, illuminance, PM10). room_id (str, req).
get_cluster_doc Perform semantic search across Matter cluster documentation (covering specifications for clusters, commands, and attributes). query (str, req); top_k (int, req).
schedule_workflow Schedule a sequential workflow of steps at a virtual absolute time. The scheduled time must be in the future relative to the current time. start_time (str, req; “YYYY-$M$-$D$ $H$:$M$:SS”); steps (list, req; e.g., {"tool":…, "args":…}).
get_current_time Get current virtual time as human-friendly string “YYYY-$M$-$D$ $H$:$M$:SS”. (none)
get_workflow_list Get list of workflows with optional filtering. (none)
A.3 List of Matter Clusters
Table 4: Implemented Matter clusters.
Cluster Attributes Commands
Basic Information VendorName, VendorID, ProductName, ProductID None
Descriptor DeviceTypeList, ServerList, ClientList, PartsList, TagList None
OnOff GlobalSceneControl, OnTime, OffWaitTime, StartUpOnOff Off, On, Toggle
Level Control CurrentLevel, RemainingTime, MinLevel, MaxLevel, CurrentFrequency, MinFrequency, MaxFrequency, OnOffTransitionTime, OnLevel, OnTransitionTime, OffTransitionTime, DefaultMoveRate, Options, StartUpCurrentLevel MoveToLevel, Move, Step, Stop, MoveToClosestFrequency
Fan Control FanMode, FanModeSequence, PercentSetting, PercentCurrent Step
MediaPlayback CurrentState Play, Pause, Stop, StartOver, Previous, Next, Rewind, FastForward
Channel ChannelList, Lineup, CurrentChannel ChangeChannel, ChangeChannelByNumber, SkipChannel
KeypadInput SupportedKeys SendKey
Identify IdentifyTime, IdentifyType Identify, TriggerEffect
Operational State PhaseList, Current Phase, CountdownTime, Operational State List, Operational State, Operational Error Pause, Resume, Stop, Start, OperationalCommandResponse
Power Source ClusterRevision, FeatureMap, Status, Order, Description, EndpointList, WiredAssessedInputVoltage, BatVoltage, BatPercentRemaining, BatChargeState, ActiveBatFaults None
Power Topology ClusterRevision, FeatureMap, AvailableEndpoints, ActiveEndpoints None
Electrical Power Measurement PowerMode, NumberOfMeasurementTypes, Accuracy, ReactiveCurrent, ApparentCurrent, ReactivePower, ApparentPower, RMSVoltage, RMSCurrent, RMSPower, Frequency, PowerFactor StartMeasurement, StopMeasurement, ResetMeasurement, GetMeasurementSnapshot
Electrical Energy Measurement Accuracy, CumulativeEnergyImported, CumulativeEnergyExported, PeriodicEnergyImported, PeriodicEnergyExported, CumulativeEnergyReset StartEnergyMeasurement, StopEnergyMeasurement, ResetCumulativeEnergy, GetEnergySnapshot
Device Energy Management ESAType, ESACanGenerate, ESAState, AbsMinPower, AbsMaxPower, PowerAdjustmentCapability, Forecast, OptOutState None
Dishwasher Mode SupportedModes, CurrentMode ChangeToMode, GetSupportedModes
Dishwasher Alarm Mask, Latch, State, Supported Reset, ModifyEnabledAlarms, GetAlarmState, GetActiveAlarms
Refrigerator And Temperature Controlled Cabinet Mode SupportedModes, CurrentMode ChangeToMode
RVC Clean Mode SupportedModes, CurrentMode ChangeToMode
RVC Operational State PhaseList, CurrentPhase, CountdownTime, Operational StateList, Operational State, OperationalError Pause, Resume, GoHome
RVC Run Mode SupportedModes, CurrentMode Start, Stop, Map, StopMap
Temperature Control TemperatureSetpoint, MinTemperature, MaxTemperature, Step, SelectedTemperatureLevel, SupportedTemperatureLevels SetTemperature
Temperature Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue None
Thermostat LocalTemperature, OccupiedCoolingSetpoint, OccupiedHeatingSetpoint, ControlSequenceOfOperation, SystemMode SetpointRaiseLower
WindowCovering Type, ConfigStatus, OperationalStatus, EndProductType, Mode, SafetyStatus, CurrentPositionLiftPercent100ths, TargetPositionLiftPercent100ths, NumberOfActuationsLift, etc. UpOrOpen, DownOrClose, StopMotion, GoToLiftPercentage
Laundry Dryer Controls SupportedDrynessLevels, SelectedDrynessLevel None
Laundry Dryer Mode SupportedModes, CurrentMode ChangeToMode
Laundry Washer Controls SpinSpeeds, SpinSpeedCurrent, NumberOfRinses, SupportedRinses None
Laundry Washer Mode SupportedModes ChangeToMode
Relative Humidity Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue, Tolerance None
A.4 List of Device Types
Table 5: List of implemented device types and their corresponding clusters.
Device type Clusters
Air Conditioner Basic Information, Fan Control, OnOff, Thermostat
Air Purifier Basic Information, Descriptor, Fan Control, Identify, OnOff
Dehumidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
Dimmable Light Basic Information, Level Control, OnOff
Dishwasher Basic Information, OnOff, Operational State
Electrical Sensor Basic Information, Electrical Energy Measurement, Electrical Power Measurement, Power Topology
Fan Basic Information, Fan Control, OnOff
Freezer Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
Heat Pump Basic Information, Descriptor, Device Energy Management, Electrical Energy Measurement, Electrical Power Measurement, Power Source, Power Topology, Thermostat
Humidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
Laundry Dryer Basic Information, Laundry Dryer Controls, Laundry Dryer Mode, OnOff, Operational State
Laundry Washer Basic Information, Laundry Washer Controls, LaundryWasherMode, OnOff, Operational State, Temperature Control
On Off Light Basic Information, OnOff
Refrigerator Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
RVC Basic Information, RVCCleanMode, RVCOperational State, RVCRunMode
TV Basic Information, Channel, KeypadInput, Level Control, MediaPlayback, OnOff
Window Covering Controller Basic Information, Window Covering
A.5 Goal Examples
Table 6: Example goals for each query types
Query Type Query Required Actions Goal
QT1 Feasible How bright is the utility room lighting right now? I am sorting some boxes and wondering if there is enough light. Also how is the living room humidity doing? I am thinking about the plants there and want to know if they are comfortable. get_room_states(utility_room) get_room_states(living_room) The utility room illuminance is 1000 lux. The living room humidity is 50%.
QT1 Infeasible I am about to shower and wondering what fan modes are available for fan 1 in the bathroom? get_room_devices(bathroom) Bathroom fan 1 not found; mode unavailable.
QT2 Feasible Ugh the kitchen feels really dry my hands are tight I left the bread rising there so I am already a bit worried about it. The living room feels dusty my eyes are itching and my throat is a little raw like there is grit in the air. get_room_devices(kitchen) get_room_devices(living_room) Increase kitchen humidity; decrease living room PM10.
QT2 Infeasible Ugh the office is so chilly, my hands go numb just thinking about working there later get_room_devices(office) Office heat pump 1 is missing; cannot increase temperature.
QT3 Feasible Set a softer light in the living room for evening reading, turn the living room dimmer light 1 on and set it to level 50. Cool the study a bit for working comfort, turn the study room AC 1 on, switch it to cooling mode and set the fan to 50 percent. get_room_devices(living_room) get_room_devices(study_room) Living room dimmable light 1 on at level 50; study room air conditioner 1 on, cooling mode, fan 50%.
QT3 Infeasible It’s a bit stuffy this morning, please turn on the bedroom air purifier 1 and set the fan to 80 percent. get_room_devices(bedroom) Not feasible: bedroom air purifier 1 is missing; cannot set fan to 80%.
QT4-1 While I am out here sorting laundry and trying to clear damp air, get the bathroom comfortable so it feels fresh by the time I walk over. Power on fan 1 in the bathroom 9 minutes from now at 30 percent, and bump it up to 40 percent 7 minutes after the prior action. Power on dimmer light 1 in the bathroom 28 minutes from now at level 10, and raise it to level 40 17 minutes after the prior action. get_room_devices(bathroom) At 9 minutes: bathroom fan 1 on, 30%. At 16 minutes: bathroom fan 1 on, 40%. At 28 minutes: bathroom dimmable light 1 on, level 10. At 45 minutes: bathroom dimmable light 1 on, level 40.
QT4-1 Temporal-Conflict Can you from the kitchen schedule dimmer light 1 in the living room to turn on and set to 80 percent in eight minutes from now, which will be 11:25 AM, I need it like that to warm up the room for guests and the start of the movie None At 8 minutes: living room dimmable light 1 on, level 80.
QT4-2 I am folding laundry and getting things ready. 20 minutes after the washer 1 in the utility room finishes, power on air purifier 1 in the living room and set the fan to 40 percent and switch heat pump 1 in the utility room to heating mode get_room_devices(living_room) get_room_devices(utility_room) At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
QT4-2 Temporal-Conflict The wash leaves the utility room humid and cool so I want the air cleaned and the space warmed right after it settles. Exactly 20 minutes after washer 1 in the utility room finishes and at 12 36 PM, turn on air purifier 1 in the living room to a gentle fan speed and turn on heat pump 1 in the utility room for heating. None At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
QT4-3 Waiting on the kitchen steam to clear so the laundry does not get musty. When dishwasher 1 in the kitchen finishes wait 11 minutes. Then start dryer 1 in the utility room. Set it to running and dryness level 1. get_room_devices(utility_room) At 99 minutes: utility room dryer 1 stopped. At 100 minutes: utility room dryer 1 running with dryness level 1.
QT4-3 Temporal-Conflict Start dryer 1 in the bathroom at twelve thirty six PM. Pause dryer 1 in the bathroom immediately when dryer 1 in the utility room finishes to avoid tripping the breaker and keep the laundry loads in order. None At 43 minutes: bathroom dryer 1 running with dryness level 1; at 44 minutes: bathroom dryer 1 paused.
A.6 LLM Judge Validation
To validate the LLM-based judging, we compared its assessments to human labels on a random subset of 70 episodes spanning all judge-scored tasks. Human annotators showed very high inter-rater reliability (Cohen’s $\kappa$ = 0.913). The LLM-Judge achieved substantial agreement with the consensus human labels (Cohen’s $\kappa$ = 0.826). These results support using the LLM-Judge as a reliable substitute for human evaluation in our benchmark.
After manually reviewing the 155 cases that the LLM-Judge evaluated as incorrect, we found that only 5 were misclassifications, underscoring the reliability of the evaluation. The detailed error distributions can be found in Table  A.11.
A.7 Experimental Setup
We evaluate 11 recent closed and open-source models: Llama-4-Scout (Meta AI, 2024), Llama-4-Maverick (Meta AI, 2024), Qwen3-32B (Yang et al., 2025), Qwen3-235B-A22B (Yang et al., 2025), Gemma-3-12B-it (Team et al., 2025), Gemma-3-27B-it (Team et al., 2025), Gemini-2.5-Flash (Comanici et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025), GPT-4.1, GPT-4.1-mini (OpenAI, 2025a), and GPT-4.1-nano (OpenAI, 2025a). All models were accessed via the OpenRouter API (OpenRouter, 2025) to ensure standardized access and comparability.
A.8 ReAct Prompt
ReAct Prompt
A.9 LLM Judge Prompt
This section presents examples of the LLM Judge Prompts that were used to evaluate smart home LLM agents. Each query type (QT) has a dedicated evaluation prompt with specific criteria.
A.9.1 QT1 Feasible Judge Prompt
QT1 Feasible Judge Prompt (Normal)
A.9.2 QT1 Infeasible Judge Prompt
QT1 Infeasible Judge Prompt
A.9.3 QT2 Infeasible Judge Prompt
QT2 Infeasible Judge Prompt
A.9.4 QT2 Infeasible-Nonexistence Judge Prompt
QT2 Infeasible Judge Prompt
A.9.5 QT3 Infeasible Judge Prompt
QT3 Infeasible Judge Prompt
A.9.6 QT4-1 Judge Prompt
QT4-1 Judge Prompt
A.9.7 QT4-2 Judge Prompt
QT4-2 Judge Prompt
A.9.8 QT4-3 Judge Prompt
QT4-3 Judge Prompt
A.10 Error Taxonomy Details
Table 7: Error Types in Feasible Episodes
Error Type Definition Example
Environment Perception Errors (EP) Failure to correctly perceive or retrieve a value of environmental variables. Querying wrong sensor, misidentifying device state, guessing instead of perceiving.
Intent Inference Errors (II) Misinterpreting user’s underlying goal. Not executing actual commands even when a user’s intention is clear.
Device Control Errors (DC) Executing the wrong device, wrong command, or missing control steps. setting wrong channel, adjusting fan speed without turning it on first.
Action Planning Errors (AP) Incorrect or incomplete construction of the control workflow. Breaking logical dependencies, only executing part of a multi-goal query without consideration.
Temporal Reasoning Errors (TR) Miscalculating relative/absolute times or sequence alignment. Scheduling “in 10 minutes” at wrong time, miscomputing dishwasher completion.
Table 8: Error Types in Infeasible Episodes
Error Type Definition Example
Contradiction Mishandling Errors (CM) The agent detects a contradiction but does not follow the proper instruction-following rule. e.g., instead of informing the user that a requested action is impossible, it arbitrarily manipulates other devices or ignores the instruction.
Contradiction Blindness Errors (CB) The agent completely fails to recognize a contradiction and executes the request as if it were valid. e.g., dimming an on/off light, scheduling conflicting temporal actions without noticing inconsistency.
LLM-Judge Errors (LJ) Errors caused not by the agent but by the evaluation system misclassifying or overlooking behavior. e.g., penalizing an informative refusal as a failure, or wrongly accepting hallucinated control as valid.
A.11 Error Type Distributions
Error Type QT2 QT3 QT4-1 QT4-2 QT4-3
Environment Perception (EP) 3 0 4 1 0
Intent Inference (II) 3 1 0 4 5
Device Control (DC) 20 7 13 13 8
Action Planning (AP) 2 0 6 3 7
Temporal Reasoning (TR) 0 0 2 6 13
Total 28 8 25 27 33
Table 9: Error type distribution of GPT-4.1 in feasible episodes.
Error Types QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
Contradiction Mishandling (CM) 8 24 6 5 6 1
Contradiction Blindness (CB) 0 5 0 40 25 30
LLM-Judge (LJ) 1 0 0 0 1 2
Total 9 29 6 45 32 33
Table 10: Error type distribution of GPT-4.1 in infeasible episodes.
A.12 Equations of Aggregators
$S_{r,t+1}=S_{r,t}+\sum_{d\in D_{S,r}}\Delta S_{d,r}(t),$ (1)
where $D_{S,r}$ denotes the set of devices in room $r$ that are defined to affect state $S$, and $\Delta S_{d,r}(t)$ represents the contribution of device $d$ at tick $t$ to $S$ in room $r$.
A.13 QT3 Error Type Distribution
Figure 6: Distribution of tool execution error types in successful QT3 Feasible cases across four models.
A.14 LLM Usage
We disclose that LLMs (OpenAI GPT-5) were used as a general-purpose assistant for language editing, including polishing wording, improving clarity, and maintaining consistency. LLMs were not involved in research design, implementation, experiments, or analysis. All scientific contributions and claims in this paper are the sole responsibility of the authors.

### v2

@@ 5.2.2 Role of Tool Feedback
5.2.3 Performance-Latency Trade-off
As discussed in §5.1, advanced reasoning models demonstrate significant performance gains. However, these gains come with a critical trade-off in latency. Table 3 shows the average episode completion time for high-performing models. Reasoning models have unacceptably large latency for practical deployment. For instance, GPT-5.1 takes up to 135.1 seconds and Gemini-2.5-Pro takes up to 74.0 seconds on average. In a smart home context, users typically expect immediate responses in less than a second. Therefore, even GPT-4.1 (best-performing standard model) cannot meet real-time requirements, as it takes up to 29.7 seconds. The extended latency of reasoning models further limits practical utility despite their accuracy improvements. At the other end of the spectrum, models under 7B parameters are more practical choices for deployment but achieved minimal performance as shown in §5.1. This suggests that the complex reasoning and tool-use coordination required by SimuHome present significant challenges for real-world smart home environments.
5.2.4 Disentangling Framework Limitations from Model Capabilities
Figure 6: Performance comparison between ReAct and HiAgent on QT4 tasks.
Table 4: Recovery rates from QT4 failures using self-review.
Type Failures Recoveries Rate Steps
QT4-1 25 2 8.0% 64.4
QT4-2 27 5 18.5% 26.8
QT4-3 33 0 0.0% 29.7
To investigate whether the failures in QT4 stem from the ReAct framework itself or from the models’ core reasoning capabilities, we conducted two ablation studies.
Adopting a Complex Planning Algorithm. We adopted HiAgent (Hu et al., 2025), a framework that incorporates hierarchical working memory and advanced planning algorithms. We compared its performance against ReAct on QT4 tasks using GPT-4.1. Figure 6 shows that HiAgent outperformed ReAct on QT4-2 and QT4-3, but underperformed on QT4-1, achieving only 40% compared to ReAct’s 50%. These results indicate that while ReAct’s framework has limited support for complex temporal coordination, not all temporal reasoning failures stem from the framework.
Enabling Runtime Periodic Self-Review. We tested a realistic setting where agents can adjust plans through periodic self-review. After scheduling workflows, the agent received callback triggers to review and correct its plans before and immediately after execution. The agent was required to independently evaluate the outcome to determine the appropriate subsequent actions. The results in Table 4 reveal that self-review achieved only 0% to 18.5% recovery rates across QT4 tasks. Crucially, even when the agent was explicitly prompted to inspect the home state immediately after execution, it failed to recognize the task failure in most cases.
The continued failure on QT4 tasks, even after adopting advanced planning techniques and enabling runtime periodic self-review, demonstrates that the primary bottleneck is not the ReAct framework itself. Instead, the failures are likely attributed to the inherently limited capabilities of current models to ground complex home states and perform precise temporal reasoning.
@@ 6 Conclusion
We propose SimuHome, a Matter-aligned simulator and benchmark that reproducibly evaluates smart home LLM agents under realistic, dynamically changing conditions. We model 4 environmental variables (i.e., temperature, illuminance, humidity, air quality) and 17 device types with time-based effects and strict reproducibility, enabling near drop-in transfer to real Matter-compliant devices. We provide 600 episodes across 12 query types with feasible and infeasible variants, packaging each episode with an initial state, a verifiable goal, a natural-language query, and required actions for process-aware, objective scoring. We score feasible tasks by final state-to-goal comparison in the simulator and assess infeasible logic checks with LLM judge that shows high agreement with human evaluation. We evaluate 16 models under the ReAct setup. Even high-performing standard models struggle with implicit intent inference, state verification, and temporal scheduling. While reasoning models consistently outperform standard models across all query types, they incur prohibitive latency costs. These results highlight a critical performance-practicality trade-off, positioning SimuHome as a vital testbed to address these challenges for real-world deployment.
Appendix A Infeasible Query Types
@@
QT1 Infeasible. This evaluates the ability to identify requests based on a false premise, such as asking for information about non-existent devices or unsupported attributes. For example, for the request “Can you tell me the vendor ID for the air purifier in the living room?”, the agent must check the list of devices in the living room, confirm the absence of an air purifier, and explain that the request’s premise is invalid.
@@
QT2 Infeasible. This assesses the ability to identify situations where, even if the user’s intent is correctly inferred, the goal is impossible to achieve due to environmental constraints or device limitations. For example, in response to “The living room feels like a sauna”, the agent must verify that the living room’s cooling system is already operating at maximum capacity and explain, with supporting reasons, why further cooling is not possible.
@@
QT3 Infeasible. This evaluates the ability to identify and reject a command to control a non-existent device. For example, for the request “Turn on the humidifier in the living room”, the agent must check the device list for the living room and confirm the absence of a humidifier. It should then explain that the request cannot be fulfilled and terminate the task without altering any device’s state.
@@
QT4-1 Infeasible. This assesses the ability to identify and explain situations where a scheduling request is invalid because the user’s specified relative and absolute times are contradictory, or because the user has a misunderstanding of the current time. For example, if a user asks, “It’s 6 p.m. now, right? Turn on the kitchen light five minutes later at 6:05 p.m.”, but the actual time is not 6 p.m., the agent must check the current time, detect the discrepancy between the relative expression “five minutes later” and the absolute time “6:05 p.m.”, and clearly explain the contradiction.
@@
QT4-2 Infeasible. This evaluates the ability to identify and explain, with evidence, requests where the user incorrectly assumes a device’s completion time or creates a contradiction by providing both relative and absolute times. For example, suppose a washer is set to finish at 6:30 p.m., but the user requests, “I think the washer finishes at 6 p.m., so start the dehumidifier at 5:50 p.m., which is 10 minutes before it finishes”. The agent must check the washer’s actual estimated completion time. It then needs to point out that the user’s assumption (6 p.m.), the relative expression (“10 minutes before”), and the absolute time (“5:50 p.m.”) are all inconsistent. The agent must not register the schedule until the contradiction is resolved and should ask the user to reconfirm the correct timing.
@@
QT4-3 Infeasible. This evaluates the ability to identify and explain that a requested deadline is physically impossible to meet, given the current progress of two operating devices. For example, if the user requests, “Guests arrive at 6 p.m., so ensure both the washer and the dishwasher are completed by 5:30 p.m.”, the agent must check the current time and the minimum time required for each device to finish. Based on this, it should explain with clear reasoning why a 5:30 p.m. completion is not feasible and suggest the earliest possible completion time or an alternative sequential plan.
Appendix B List of Tools
Tool List for Agent
@@ Appendix B List of Tools
[…]
Name Description Args
[…]